Calculate the hydroxide ion concentration at 25 in a solution whose pH is 5.78. (kW=1*10_14moldm at25)


1
Expert's answer
2022-03-23T14:09:06-0400

Solution:

hydroxide ion = [OH−]


pH = −log[H3O+]

Therefore,

[H3O+] = 10−pH = 10−5.78 = 1.66×10−6

[H3O+] = 1.66×10−6 M


Kw = [H3O+] × [OH−]

Therefore,

[OH−] = Kw / [H3O+] = (1.0×10−14) / (1.66×10−6) = 6.02×10−9

[OH−] = 6.02×10−9 M


Answer: The hydroxide ion (OH−) concentration is 6.02×10−9 M